The speed of a stream is 5mph. A boat travels 5 miles upstream in the same time it takes to travel 15 miles downstream. What is the speed of the boat in still water.



help please.

Answer by mananth(16949)

boat speed x mph
current speed 5 mph
..
against current x - 5 mph
with current x + 5 mph

Distance against current 5
Distance with current 15
Time = same
5 /(x- 3 )= 15 /(x+ 3 )
5 *(x+ 3 )= 15 *(x- 3 )
5 x + 15 = 15 x -45
15 x -5 x = 45 + 15
10 x = 60
/ 10
x= 6 mph